• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Danagoulian

    Log entry time 00:36:52 on November16,2001

    Entry number 77311

    keyword=shift summary

    Swing shift
    SL : A. Danagoulian
    shift worker : A. Camsonne

    18:15 run 2177 last I+ run ended
    configuration change from I+ to q measurement
    moved left spectrometer collimator to sieve slits
    starting moving right spectrometer from 84.17 to 56.22 degree
    vernier cameras on both spectrometers don't move any more

    18:51 right spectrometer at 56.22 degrees
    changed prescale factor to 500/30/51/35/1

    19:33 Controlled access to fix the vernier cameras of the spectrometers

    19:45 started moving right spectrometer to 32.95 degrees

    20:20 Finished moving spectrometer and collimator to OPEN
    Beam back
    Changed prescale factor to 2290/100/1530/168/1

    20:48 Changed prescale factor to 4580/200/3060/336/1

    21:39 Coda problems
    Roc 14 crashed
    killed coda and rebooted ROC 14
    following runs tend to have 100 % dead time
    killed CODA and rebooted all ROCs, it seems that it fixed the problem for a good amount of time


    22:00-24:00 Taking data.